OpenStudy (anonymous):

To dilute a concentrated liquid fabric softener, the directions state to mix 3 cups of water with 1 cup of concentrated liquid. How many gallons of water will you need to make 6 gallons of diluted fabric softener?

OpenStudy (anonymous):

easy, line it up: 3/1=(how many cups are in a gallonx6)/x

OpenStudy (anonymous):

sooo 3/1=16/x ?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

erm no what is 16x6?

OpenStudy (anonymous):

96 lolol

OpenStudy (anonymous):

alright so 96/x=3/1

OpenStudy (anonymous):

96=3x

OpenStudy (anonymous):

and then 96/3=x

OpenStudy (anonymous):

so 32=x

OpenStudy (anonymous):

ohhh okay so would the answer be 2 gallons of water?

OpenStudy (anonymous):

yea i guess... is it a multiple choice?

OpenStudy (anonymous):

yes lolol a. 1 1/2 gallons b. 3 gallons c. 2 1/2 gallons d. 4 gallons e. 4 1/2 gallons

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

try 96/x=1/3?

OpenStudy (anonymous):

idk man im tired and i got a lot of work... im srry

OpenStudy (anonymous):

hahah its okay